The units digit of a two-digit number is twice the tens digit. 1/2 the sum of the digits is equal to 2 less than the units digit. What is the number?
Answer by ReadingBoosters(3246)

x = units
y = tens
x = 2y
(1/2)(x + y) = x - 2
Substitute x into the second equation
(1/2)(2y + y) = 2y - 2
Combine like terms
(1/2)(3y) = 2y - 2
1.5y = 2y - 2
1.5y - 2y = 2y - 2 - 2y
-.5y = -2
-.5y/-.5 = -2/-.5
y = 4
Solve x using y
x = 2y = 2(4) = 8
x is the unit 8
y is the digit 4
The number is 48